# Finiteness of leaps of modules of integrable derivations of algebras of finite type

### v1

@@ 4 The local case
satisfies the required properties (note that $l_{m+1}=m^{\prime}p^{i-\tau}$). This completes the proof in the case $i\geq\tau$.
@@
satisfies the required properties. This completes the proof in the case $i<\tau$. ∎
@@
satisfy the required properties. This completes the proof. ∎
@@
where the $C_{i}$ are as in 4.2-4.6. Let $\operatorname{Der}_{k}^{q}(R,R)$ be the module of $q$-integrable derivations of $R$. Then any element in $\mathfrak{n}^{e(q+1)+1}\cdot\operatorname{Der}_{k}^{q}(R,R)$ is $m$-integrable for any positive integer $m$. In particular, the set of leaps produced by $\operatorname{Der}_{k}^{q}(R,R)$ is finite.
@@
Suppose that $d\in\operatorname{Der}_{k}^{q}(R,R)$ and $x\in\mathfrak{n}^{e(q+1)+1}$. By assumption, there exists $\widetilde{E}\in\operatorname{HS}_{k}^{q}(R)$ such that $\widetilde{E}_{1}=d$. We can lift $\widetilde{E}$ to $E\in\operatorname{HS}_{k}^{\infty}(T)$ by formal smoothness of $T$. Let $y\in\mathfrak{m}^{e(q+1)+1}$ be a preimage of $x$. Then $D:=y\bullet E$ is $I$-logarithmic up to $q$ and logarithmically $a[e,q,q]$-bounded (note that $a[e,q,q]=(e(q+1),2e(q+1),3e(q+1),\dots)$). Let $m$ be an arbitrary positive integer. Applying the above proposition recursively, we can find $D^{+}\in\operatorname{HS}_{k}^{\infty}(T)$ such that $D^{+}$ is $I$-logarithmic up to $m$ and $D$ and $D^{+}$ coincide up to $\min\{l_{q+1}-1,l_{q+2}-1,\dots,l_{m}-1\}.$ Note that $l_{n}\geq 2$ for any $n>q$. In particular, $y\cdot E_{1}=D^{+}_{1}$. This means that $x\cdot d$ is $m$-integrable. This shows the first part of the theorem. The second part is a consequence of the following theorem. ∎
The next theorem is a slight variant of [bravo2024finitenessleapssensehasseschmidt, Theorem 3.1], [NARVAEZMACARRO2024109441, Proposition 1.11] and its proof is essentially the same. For the convenience of the reader, we include the proof here as well.
@@
$a_{1}\overline{d}_{s_{1}}+\dots+a_{n}\overline{d}_{s_{n}}=0,\ a_{r}\in K$
@@
$b_{1}d_{s_{1}}=-b_{2}d_{s_{2}}-\dots-b_{n}d_{s_{n}}+d,\ b_{r}\in R$
where $b_{r}$ are lifts of $a_{r}$ to $R$ and where $d\in\mathfrak{l}M$. Note that $b_{1}$ is invertible, so that
@@
By assumption, there exists $D_{r}\in\operatorname{HS}_{k}^{s_{r}}(R)$ whose first component is equal to $d_{s_{r}}$ for each $r$. In addition, there exists $D\in\operatorname{HS}_{k}^{t}(R)$ whose first component is equal to $d$ with $t\geq s_{2}$. Then the first component of
$b_{1}^{-1}\bullet(((-b_{2})\bullet D_{2})\circ\dots\circ((-b_{r})\bullet D_{r})\circ D)\in\operatorname{HS}_{k}^{s_{2}}(R)$
equals to $d_{s_{1}}$. This means that $d_{s_{1}}$ is at least $s_{2}$-integrable, which contradicts the choice of $d_{s_{1}}$. This proves the claim, and the theorem follows from it. ∎
@@
$\{p,p^{2},\dots,q\}\cup\operatorname{Leaps}(\operatorname{Der}_{k}^{q}(R,R)),$
@@
5 The global case
In this section, we apply the results of the previous section to prove finiteness-of-leaps for arbitrary $k$-schemes essentially of finite type.
@@ 5 The global case
We construct inductively a sequence of pairs $\{(F_{n},i_{n})\}_{n=1,2,\dots}$ such that:
@@
    $F_{1}\supset F_{2}\supset F_{3}\supset\dots$
@@
i.e., $F_{\tau}=\operatorname{Ob}_{X}^{p^{i_{\tau}}}$. By (iii), $X$ satisfies $\operatorname{FL}(\tau)$ and in particular the set of leaps of $\operatorname{Der}_{k}(\mathcal{O}_{X})$ is finite.
@@
where $\iota:Y\to X$. In other words, we have $F_{n+1}=F_{n}\times_{\iota_{*}\iota^{*}F_{n}}\iota_{*}\operatorname{Ob}_{Y}^{i_{n+1}}$. This implies quasi-coherence, and hence coherence of $F_{n+1}$ in the sense stated in (ii). Condition (iii) follows from the fact that $Y$ satisfies $\operatorname{FL}(i_{n+1})$ and from the construction. The restriction of the above diagram to $Y$ is
@@
Let $R$ be a $k$-algebra essentially of finite type, $m,\tau\geq 1$ integers and $q:=p^{\tau}$. Suppose that $R$ satisfies $\operatorname{FL}(\tau)$. Let $D\in\operatorname{HS}_{k}^{m}(R)$. Let $l_{1},l_{2},\dots$ be the sequence of integers defined as in 4.11. Then there exists $D^{+}\in\operatorname{HS}_{k}^{m+1}(R)$ such that $D$ and $D^{+}$ coincide up to $l_{m+1}-1$, i.e., $D_{1}=D^{+}_{1},\dots,D_{l_{m+1}-1}=D_{l_{m+1}-1}^{+}$.
@@
Let $r:={m+1}/l_{m+1}$. As $\operatorname{Ob}_{R}^{p^{\tau}}=\operatorname{Ob}_{R}^{p^{\tau+1}}=\dots$, there exists $E\in\operatorname{HS}_{k}^{r-1}(R)$ such that $\operatorname{ob}_{m+1}(D)=-\operatorname{ob}_{r}(E)$. Then we can take $D^{+}:=D\circ E[l_{m+1}].$
@@
Let $R$ be a $k$-algebra essentially of finite type, $n\geq 1$ be an integer and $D\in\operatorname{HS}_{k}^{n}(R)$. Then $D$ is $\infty$-integrable if and only if $D$ is $m$-integrable for every integer $m\geq n$. Here, we say that $D$ is $m$-integrable if $D$ is in the image of $\operatorname{HS}_{k}^{m}(R)\to\operatorname{HS}_{k}^{n}(R)$.
@@
is a well-defined Hasse-Schmidt derivation. Note that $E^{\infty}$ restricts to $D$ by the choice of $N$. Thus, $D$ is $\infty$-integrable. ∎
@@
$\operatorname{Der}_{k}^{M}(R)=\operatorname{Der}_{k}^{M+1}(R)=\dots=\operatorname{Der}_{k}^{\infty}(R).$
In particular, $\operatorname{Der}_{k}^{\infty}(R)$ is compatible with localization of $R$.
@@
Theorems 5.1 and 5.2 hold also in the case where the base field is perfect.
@@
This is a consequence of [hernandez2020behavior, Theorem 3.27]. ∎

### v4

@@ 4. The local case
satisfies the required properties: We have just shown that $E$ is logarithmically $b[e,m+1,q]$-bounded. As each $\widetilde{E}^{j}$ is $I$-logarithmic up to q1q-1, each $F^{j}$ is $I$-logarithmic up to $(q-1)m^{\prime}p^{i-\tau}+m^{\prime}p^{i-\tau}-1=m$. Hence, also $E$ is $I$-logarithmic up to $m$ and induces
$\widetilde{E}=\widetilde{D}\circ(x_{1}\bullet\widetilde{D}^{1})[m^{\prime}p^{i-\tau}]\circ\dots\circ(x_{n}\bullet\widetilde{D}^{n})[m^{\prime}p^{i-\tau}]\in\operatorname{HS}^{m}_{k}(R).$
From this expression, we obtain $\operatorname{ob}_{m}(\widetilde{E})=0$. Finally, note that for each $j$, the first $m^{\prime}p^{i-\tau}-1$ terms of $F^{j}$ are zero and note that $l_{m+1}=m^{\prime}p^{i-\tau}$. This implies that $D$ and $E$ coincide up to $l_{m+1}-1$. This completes the proof in the case $i\geq\tau$.
@@
satisfies the required properties for the same reasons as before. This completes the proof in the case $i<\tau$. ∎
@@
satisfy the required properties: (ii) and (iii) follow from the construction. For (i), we note that
$D^{+}_{m+1}=E_{m+1}+x_{1}d_{1}+\dots+x_{n}d_{n}.$
This induces $f-f=0\in\operatorname{Hom}_{R}(I/I^{2},R)$ so that $D^{+}$ is $I$-logarithmic up to $m+1$. This completes the proof. ∎
@@
where the $C_{i}$ are as in 4.2-4.6. Let $\operatorname{Der}_{k}^{q}(R)$ be the module of $q$-integrable derivations of $R$. Then any element of $\mathfrak{n}^{e(q+1)+1}\cdot\operatorname{Der}_{k}^{q}(R)$ is $m$-integrable for any positive integer $m$. In particular, the set of leaps produced by $\operatorname{Der}_{k}^{q}(R)$ is finite.
@@
Suppose that $d\in\operatorname{Der}_{k}^{q}(R)$ and $x\in\mathfrak{n}^{e(q+1)+1}$. By assumption, there exists $\widetilde{E}\in\operatorname{HS}_{k}^{q}(R)$ such that $\widetilde{E}_{1}=d$. We can lift $\widetilde{E}$ to $E\in\operatorname{HS}_{k}^{\infty}(T)$ by formal smoothness of $T$. Let $y\in\mathfrak{m}^{e(q+1)+1}$ be a preimage of $x$. Then $D:=y\bullet E$ is $I$-logarithmic up to $q$ and logarithmically $a[e,q,q]$-bounded (note that $a[e,q,q]=(e(q+1),2e(q+1),3e(q+1),\dots)$). Let $m$ be an arbitrary positive integer. Applying the above proposition recursively, we obtain $D^{+}\in\operatorname{HS}_{k}^{\infty}(T)$ such that $D^{+}$ is $I$-logarithmic up to $m$ and $D$ and $D^{+}$ coincide up to $\min\{l_{q+1}-1,l_{q+2}-1,\dots,l_{m}-1\}.$ Note that $l_{n}\geq 2$ for any $n>q$. In particular, $y\cdot E_{1}=D^{+}_{1}$. This means that $x\cdot d$ is $m$-integrable. This shows the first part of the theorem. The second part is a consequence of the following theorem. ∎
The next theorem is a slight variant of [1, Theorem 3.1], [10, Proposition 1.11] and its proof is essentially the same. For the convenience of the reader, we include the proof here as well.
@@
$a_{1}\overline{d}_{s_{1}}+\dots+a_{n}\overline{d}_{s_{n}}=0,\ a_{i}\in K$
@@
$b_{1}d_{s_{1}}=-b_{2}d_{s_{2}}-\dots-b_{n}d_{s_{n}}+d,\ b_{i}\in R$
where $b_{i}$ are lifts of $a_{i}$ to $R$ and where $d\in\mathfrak{l}M$. Note that $b_{1}$ is invertible, so that
@@
By assumption, there exists $D_{i}\in\operatorname{HS}_{k}^{s_{i}}(R)$ whose first component is equal to $d_{s_{i}}$ for each $i$. In addition, there exists $D\in\operatorname{HS}_{k}^{t}(R)$ whose first component is equal to $d$ with $t\geq s_{2}$. Then the first component of
$b_{1}^{-1}\bullet(((-b_{2})\bullet D_{2})\circ\dots\circ((-b_{n})\bullet D_{n})\circ D)\in\operatorname{HS}_{k}^{s_{2}}(R)$
is equal to $d_{s_{1}}$. In the above expression, $D_{i}$ and $D$ are understood to be truncated to length $s_{2}$. This implies that $d_{s_{1}}$ is at least $s_{2}$-integrable, which contradicts the choice of $d_{s_{1}}$. This proves the claim, and the theorem follows from it. ∎
@@
$\{p,p^{2},\dots,q\}\cup\operatorname{Leaps}(\operatorname{Der}_{k}^{q}(R)),$
@@
5. The global case
In the previous section, we proved that if a local $k$-algebra $(R,\mathfrak{n})$ essentially of finite type satisfies the finiteness of leaps except the closed point, then so does $R$ itself (Setting 4.1, Theorem 4.17). In this section, we apply this result to prove the finiteness of leaps for arbitrary $k$-schemes essentially of finite type.
@@ 5. The global case
We construct inductively a sequence of pairs $(F_{1},i_{1}),(F_{2},i_{2}),\dots$ such that:
@@
    $F_{1}\supset F_{2}\supset F_{3}\supset\dots$,
@@
i.e., $F_{\tau}=\operatorname{Ob}_{X}^{p^{i_{\tau}}}$. By (iii), $X$ satisfies $\operatorname{FL}(\tau)$ and in particular the set of leaps of $X$ is finite.
@@
where $\iota:Y\to X$. In other words, we have $F_{n+1}=F_{n}\times_{\iota_{*}\iota^{*}F_{n}}\iota_{*}\operatorname{Ob}_{Y}^{i_{n+1}}$. This implies the quasi-coherence, and hence the coherence of $F_{n+1}$ in the sense stated in (ii). Condition (iii) follows from the fact that $Y$ satisfies $\operatorname{FL}(i_{n+1})$ and from the construction. The restriction of the above diagram to $Y$ is
@@
As an affine case of the above theorem, we obtain the following.
@@
We now study $\infty$-integrability based on the finiteness of leaps. We begin with the following lemma.
@@
Let $R$ be a $k$-algebra essentially of finite type, let $m,\tau\geq 1$ be integers, and let $q:=p^{\tau}$. Suppose that $R$ satisfies $\operatorname{FL}(\tau)$. Let $D\in\operatorname{HS}_{k}^{m}(R)$. Let $l_{1},l_{2},\dots$ be the sequence of integers defined as in 4.11. Then there exists $D^{+}\in\operatorname{HS}_{k}^{m+1}(R)$ such that $D$ and $D^{+}$ coincide up to $l_{m+1}-1$, i.e., $D_{1}=D^{+}_{1},\dots,D_{l_{m+1}-1}=D_{l_{m+1}-1}^{+}$.
@@
Let $r:={m+1}/l_{m+1}$. As $\operatorname{Ob}_{R}^{p^{\tau}}=\operatorname{Ob}_{R}^{p^{\tau+1}}=\dots$, there exists $E\in\operatorname{HS}_{k}^{r-1}(R)$ such that $\operatorname{ob}_{m+1}(D)=-\operatorname{ob}_{r}(E)$. Then $D\circ E[l_{m+1}]$ extends to $D^{+}\in\operatorname{HS}_{k}^{m+1}(R)$, whose first $l_{m+1}-1$ terms coincide with those of $D$. ∎
@@
Let $R$ be a $k$-algebra essentially of finite type, let $n\geq 1$ be an integer, and let $D\in\operatorname{HS}_{k}^{n}(R)$. Then $D$ is $\infty$-integrable if and only if $D$ is $m$-integrable for every integer $m\geq n$. Here, we say that $D$ is $m$-integrable if $D$ is in the image of $\operatorname{HS}_{k}^{m}(R)\to\operatorname{HS}_{k}^{n}(R)$.
@@
is a well-defined HS-derivation. Note that $E^{\infty}$ restricts to $D$ by the choice of $N$. Thus, $D$ is $\infty$-integrable. ∎
Setting $n=1$ in the above theorem, we obtain the following.
@@
$\operatorname{Der}_{k}^{M}(R)=\operatorname{Der}_{k}^{M+1}(R)=\dots=\operatorname{Der}_{k}^{\infty}(R)=\operatorname{Ider}_{k}(R).$
Corollary 5.6.
Let $R$ be a $k$-algebra essentially of finite type. Then $\operatorname{Der}_{k}^{\infty}(R)=\operatorname{Ider}_{k}(R)$ is compatible with localization of $R$. That is, we have
$\operatorname{Ider}_{k}(S^{-1}R)=S^{-1}\operatorname{Ider}_{k}(R)$
as $S^{-1}R$-modules for any multiplicatively closed subset $S\subset R$.
Proof.
This is a consequence of Theorem 2.3. ∎
Thus, $\operatorname{Ider}_{k}(R)$ constitutes a coherent sheaf $\operatorname{Ider}_{k}(\mathcal{O}_{X})$ on a $k$-scheme $X$ essentially of finite type. Using this sheaf, we can state Theorem 5.5 in terms of schemes.
@@
Let $X$ be a $k$-scheme essentially of finite type. Then there exists $M>0$ such that
$\operatorname{Der}_{k}^{M}(\mathcal{O}_{X})=\operatorname{Der}_{k}^{M+1}(\mathcal{O}_{X})=\dots=\operatorname{Ider}_{k}(\mathcal{O}_{X}).$
Finally, we conclude the paper by noting that the above results extend to the case of perfect fields.
Theorem 5.8.
Let $k^{\prime}$ be a perfect field of positive characteristic and let $R$ be a $k^{\prime}$-algebra of finite type. Then the set of leaps of $R$ as a $k^{\prime}$-algebra is finite.
@@
This follows from Theorem 5.2 and the fact that leaps are stable under separable base changes [3, Theorem 3.27]. ∎
Acknowledgements
The author would like to thank Professor Keiji Oguiso and the members of Oguiso’s laboratory for their interest in this work and their valuable comments and suggestions.